# Distributed acoustic sensing for ocean applications

### v2

@@ IV.1 Whale calls
where $f_{a}$ is estimated by the temporal gradient of the phase (29). Figure 10 shows (a) the acoustic frequency and (b) the incidence angle as estimated through the temporal and spatial gradient of the phase of the complex-valued data, respectively.
To highlight the characteristics of the V-shaped arrivals attributed to whale vocalizations, Fig. 11 details the phase (28) and the spatial phase gradient (30) of the earliest arrival, recorded at around 10 s, and compares the results with a fitted simulation example. The simulated wave is a sinusoidal pulse of 1 s duration, at a frequency of 18 Hz as estimated from the temporal phase gradient (29). The simulated pulse is propagating at a sound speed of 1500 m/s and the point of closest approach is set at $z_{\rm pca}=28.4$ km as inferred from the apex of the wavefront arrival in Fig. 8(c). Considering plane wave incidence at an angle of 70 as estimated from  (30), the propagation delay along the $z$ axis is calculated as $t=(z-z_{\rm pca})\sin\theta/c$. With the plane wave assumption, the simulation in Figs. 11(c) and (d) is already a good approximation of the recorded data, especially at distances far from the point of closest approach. Considering spherical, omnidirectional wave propagation from a source at a range of $r=1$ km from the point of closest point of approach the propagation delay along the $z$ axis is $t=(\sqrt{(z-z_{\rm pca})^{2}+r^{2}}-r)/c$. In this case, the simulation in Figs. 11(e) and (f) approximates the data more accurately also at distances around the point of closest approach. Figure 12 shows the wavefront angle of incidence, $\theta=\arctan\left((z-z_{\rm pca})/r\right)$, as a function of the distance from the point of closest approach $z-z_{\rm pca}$ for different source ranges $r$.
The uncertainty in the channel locations, the sound speed profile, and the low signal-to-noise ratio make source localization by backpropagation of the amplitude or phase of the recorded signal challenging. Backpropagating the spatial phase gradient $\Delta_{z}\phi$ of the recorded signal allows an accurate estimation of the location along the cable where the source has its closest point of approach (CPA). Consider a two-dimensional grid in time $t_{0}\equiv t$ and space $z_{0}\equiv z$ with potential spatiotemporal source locations $(t_{0},z_{0})$, i.e., each point on the grid represents a potential apex of a wavefront as shown in Fig. 11. Assuming plane wave propagation at $\theta=90$, a reasonable approximation for $z-z_{\rm pca}>10$ km as shown in Fig. 12, summing the data points over all channels $z$ (that is channels 10000–20000) along the wavefront delays $\left|z-z_{0}\right|/c_{a})$ populates the $(z_{0},t_{0})$ grid as,
@@
The two-dimensional backpropagation result normalized to the maximum absolute value $\mathbf{B}/\max(\left|\mathbf{B}\right|)$ is shown in Fig. 13(a). Since the spatial phase gradient changes sign symmetrically around the wavefront’s normal incidence, it will sum up most destructively at the point of closest approach. Figure 13(b) shows the spatial covariance matrix $\mathbf{C}_{b}=\mathbf{B}\mathbf{B}^{T}$ of the two-dimensional backpropagation output $\mathbf{B}$ in Fig. 13(a). The point of closest approach is identified by locating the minimum of the diagonal of the spatial covariance matrix, $z_{\rm pca}=\arg\min\left(\text{diag}(\mathbf{C}_{b})\right)$; see Fig. 13(c).
Figure 14: (color online) High-pass filtered (cut-off frequency 15 Hz) data featuring a broadband signal attributed to a passing ship. (a) Recordings along successive channels in the north cable optical fiber. Two faint fin whale calls can be observed at around 10 and 20 seconds into the recording. (b) Data representation in the f-k domain. The frequency sweep, centered around 20 Hz, is attributed to the fin whale calls which overlap with the ship’s signal.
Figure 15: (color online) Estimates of (a) the frequency $f_{a}$ (29) and (b) the angle of incidence $\theta$ (30) of the acoustic wavefield sensed by the optical fiber cable.
@@ IV.2 Ship noise
Dataset B was collected along the transmit fiber in the north cable on 3/11/2021. The duration of the recording is 30 s sampled at 500 Hz. The channel spacing is $L_{h}=2.0419$ m and the gauge length is set to $L_{g}=15L_{h}\approx L_{p}\approx 30$ m.
Similar to Fig. 8, Fig. 14 shows the time-space and f-k representation of the signal recorded over 3000 channels (channels 22500–25500) after de-trending, transforming the measurements to strain units and high-pass filtering at a cut-off frequency of 15 Hz. The recorded signal in dataset B indicates omnidirectional excitation in the frequency range 40–60 Hz, which is attributed to ship noise.wilcock2023distributed The notches along the wavenumber dimension occur at multiples of $k=\pm 1/L_{g}\approx\pm 0.033$ m-1.
Figure 15 shows (a) the acoustic frequency and (b) the incidence angle as estimated through the temporal and spatial gradient of the phase of the complex-valued data, respectively. Finally, Fig. 16 shows the result of backpropagating the spatial phase gradient $\Delta_{z}\phi$ of the recorded signal and the estimation of the closest point of approach, similar to Fig. 13.
Figure 17: (color online) T wave signal from the Mw 5.2 Fox Islands earthquake on November 4, 2021 at 08:57:06 UTC. (a) Map with great circle arrival of T-wave, (b) bathymetry and time-space figure recorded on the north cable. (c) waveform and spectrogram of a receiver at 40 km.
@@ IV.3 Earthquake noise
T waves, i.e., acoustic waves coupled into the ocean at a sloping seafloor interface, such as a seamount or the continental margin, due to seismic body waves from nearby earthquakes,okal2008generation can be observed with underwater DAS cables at large epicentral distances.Ugalde2022 T waves from several earthquakes have been observed in the OOI DAS data.shen2024ocean Among these, the Mw 5.2 Fox Islands earthquake, which occurred on November 4, 2021, at 08:57:06 UTC, is one of the largest events recorded during this experiment.shen2024ocean
@@
The T wave is observed around 30 km along the cable, which can be due to a change of the bathymetry reducing the modal cut-off frequency. The reduction of the T wave energy at approximately 60 km along the cable is attributed to the directional sensitivity of DAS, explained in Sec. II.5, since the T wave signal arrives at a 120 heading, resulting in a horizontal incident angle of 60–70 with the cable; see Fig. 17(a).
@@ V.1 Optical multiplexing
Most DAS applications have used dark fibers, i.e., redundant network fibers not used for communication, to avoid interference, as overlapping DAS pulses with existing optical traffic can degrade both communication and DAS data quality. Most fibers in subsea networks are actively used for telecommunications, unlike terrestrial networks where dark fibers are more readily available due to lower relative installation costs. Consequently, dark subsea fibers are rare, and the high demand for intercontinental communication further limits their availability. Optical multiplexing methods based on frequency division can overcome this limitation.
Ordinary infrared optical communication typically occurs in the C-band (“Communication”), 1530–1565 nm, divided into channels. The information to be transmitted is optically encoded by a transceiver operating at one channel’s wavelength (or equivalently, frequency), and the output of all transceivers is combined into the same fiber—a technique known as wavelength domain multiplexing (WDM)—to permit simultaneous transmission of independent information. Initially, simultaneous operation of DAS and communication traffic in the same fiber was demonstrated with C-band WDM, dedicating three channels for the DAS.huang2019first ; jia2021experimental ; guerrier2022vibration The disadvantages of this approach are that the communication bandwidth is reduced relative to a fiber without DAS and the peak DAS power must be limited to prevent interference effects.jia2021experimental
Another approach utilizes WDM of C-band communications with DAS operating in the L-band (“Long”), 1565–-1625 nm, which has only slightly higher attenuation than the C-band but is used much less frequently in communications.wienecke2023new L-band multiplexed DAS has been demonstrated both in laboratory,wienecke2023new and in field experiments.brenne2024non ; shi2024multiplexed
@@ V.2 Measurement range
Commercially available DAS systems have a measurement range up to 170 km.waagaard2021realtime Most subsea cables utilize repeaters at regular intervals every 40–100 km to amplify the optical signal and counteract transmission loss. However, common repeaters are effectively one-way gates preventing any backscattered signal from returning back to the interrogator. Consequently, the DAS measurement range is limited to the fiber length that spans the distance between the interrogator and the successive repeater.
@@ V.3 High-frequency signal observation
While most studies have demonstrated the capabilities of DAS in observing hydro-acoustic signals at frequencies below 100 Hz—such as those produced by baleen whales and ship noise—the maximum frequency for DAS systems is inherently determined by the pulse repetition rate of DAS recordings, which is limited by the two-way travel time of optical pulses along the cable. For instance, a 50 km cable allows a pulse repetition frequency of $f_{r}=c_{n}/(2L_{f})$=2 kHz resulting in a maximum detectable frequency of $f_{\text{max}}=f_{r}/2=1$ kHz. However, several factors can reduce DAS sensitivity to high-frequency signals:
@@
Despite these challenges, several case studies have demonstrated the potential of DAS for observing acoustic signals above 100 Hz. Specifically, ship noise was observed with DAS up to 120 HzThiem2023ShipNC ; cao2023near and a sound source emitting frequencies up to 160 Hz was utilized to localize a DAS cable in shallow watershen2024high , whereas Ref. taweesintananon2021distributed, demonstrated DAS observing seismic airgun pulses up to 960 Hz. In another study,douglass2023distributed a 3.5 km subsea cable sampled at 2 kHz captured impulsive signals from a low-power acoustic source near the surface, measuring frequencies up to 700 Hz. A recent study in Lake Zurich, Switzerland, used a 1 km custom cable designed for DAS to record an acoustic source at frequencies up to 2500 Hz,rychen2023test a similar maximum frequency as demonstrated for communication signals.potter2024first
@@ V.4 Unknown geometry
The exact channel locations in DAS are often unknown, unlike hydrophone arrays in ocean acoustics. Particularly, the precise positioning for most telecom cables is either unavailable or proprietary information. Even when the overall cable location is known, estimating the precise channel locations remains challenging. The exact speed of light in optical fiber cannot be measured in situ, leading to uncertainty in converting fast time to distance. Previous studies have demonstrated that a sound source broadcasting near the surface can localize the cable in shallow water. shen2024high By extension, ships, when combined with AIS location data, can serve as opportunistic sources to help determine the cable’s position.
@@ VI Ocean applications beyond acoustics
Since fiber-optic cables deployed on the seafloor are subject to a wide range of perturbations, including strain, temperature, and pressure originating from ocean waves, bottom currents, earthquakes, and slope movement, this section presents applications of distributed acoustic sensing in oceanography and seismology, beyond acoustics.
@@ VI.1 Ocean-bottom seismology
Figure 18: (color online) (a) Ambient seismic noise (Scholte waves) along the OOI RCA North cable between channels 9000–14000 (depth ca. 200 m), band-pass filtered in the frequency range 0.2–1 Hz. (b) Frequency-wavenumber spectrum of the same data segment. Positive wavenumbers indicate propagation towards the shore. (c), (d) Same as (a), (b) except for channels 27500–32500 (depth ca. 500 m) in deeper water.
The paucity of long-term seismographic stations in the oceans is a challenge in global seismology, inhibiting earthquake monitoring and early warning for offshore fault zones and limiting ray path coverage across broad regions of the core and mantle in tomographic models.kohler2020plan Fiber-optic sensing on subsea cable networks offers a promising solution to increase global coverage. DAS recordings of earthquakes on subsea cables have been utilized for event detection and location,lior2021detection ; baba2023observation identification of offshore fault zones from scattered waves,lindsey2019illuminating and quantification of linear and nonlinear site effects for strong ground motion.viens2022nonlinear Because the DAS measurement is made onshore in the interrogator, onshore-offshore telemetry occurs at the speed of light. Recent tests integrating ocean-bottom DAS with real-time earthquake early warning, suggest a significant reduction in warning time (up to 10 s) for earthquakes nucleating on offshore fault zones.romanowicz2023seafoam ; yin2023real
Ocean-bottom DAS offers a unique opportunity to probe the shear-wave velocity structure of marine sediments and oceanic crust, with good horizontal and frequency ($<$10 Hz) resolution. DAS can outperform conventional ocean-bottom seismometers as horizontal-component seismic sensors,janiszewski2023broad since the low profile of cables above the seafloor reduces the influence of ocean currents and the compliance noise from ocean waves can be removed by array processing (see below). The horizontal orientation makes DAS sensitive to shear waves from earthquakes and shear-wave resonances in seismic ambient noise, which can be used to infer the thickness of submarine sediments.trabattoni2024sediment ; taweesintananon2024near
For plane harmonic waves, the axial strain $\epsilon_{zz}$ can be expressed as the ratio between the axial particle velocity $v_{z}$ and the apparent phase velocity $c=\omega/(k_{a}\sin\theta)$
$\epsilon_{zz}=\frac{\partial u_{z}}{\partial z}=\frac{\partial u_{z}}{\partial t}\frac{\partial t}{\partial z}=\frac{v_{z}}{c},$ (32)
which is equivalent to (19).wang2018ground Hence, ocean-bottom DAS is particularly sensitive to low-phase velocity or short-wavelength phases, like Scholte waves (water-sediment interface modes). Array analysis of Scholte waves with meter-to-kilometer scale resolution permits recovery of frequency-dependent multi-modal dispersion (Fig. 18), which has been widely utilized to invert for the depth-dependent shear-wave velocity structure of shallow marine sediments. spica2020marine ; cheng2021utilizing ; williams2021scholte Ocean-bottom DAS observations of seismic noise—for example, the 0.2–1 Hz noise radiating from the coast evident in f-k spectra from the OOI RCA (Fig. 18)—has been used to characterize microseisms.williams2019distributed ; guerin2022quantifying ; xiao2022locating ; fang2022unraveling

### v3

@@ IV.1 Whale calls
where $f_{a}$ is estimated by the temporal gradient of the phase, Eq. (35). Figure 11 shows (a) the acoustic frequency and (b) the incident angle as estimated through the temporal and spatial gradient of the phase of the complex-valued data, respectively.
To highlight the characteristics of the V-shaped arrivals attributed to whale vocalizations, Fig. 12 details the phase, Eq. (34), and the spatial phase gradient, Eq. (36), of the earliest arrival recorded at around 10 s, and compares the results with a fitted simulation example. The simulated wave is a sinusoidal pulse of 1 s duration, at a frequency of 18 Hz as estimated from the temporal phase gradient, Eq. (35). The simulated pulse is propagating at a sound speed of 1500 m/s and the closest point of approach (CPA) is set at $z_{\rm cpa}=28.4$ km as inferred from the apex of the wavefront arrival in Fig. 9(c). Considering plane wave incidence at an angle of 70 as estimated from Eq. (36), the propagation delay along the $z$ axis is calculated as $t=(z-z_{\rm cpa})\sin\theta/c$. With the plane wave assumption, the simulation in Figs. 12(c) and (d) is already a good approximation of the recorded data, especially at distances far from the point of closest approach. Considering spherical, omnidirectional wave propagation from a source at a range of $r=1$ km from the closest point of approach the propagation delay along the $z$ axis is $t=(\sqrt{(z-z_{\rm cpa})^{2}+r^{2}}-r)/c$. In this case, the simulation in Figs. 12(e) and (f) approximates the data more accurately also at distances around the closest point of approach. Figure 13 shows the wavefront angle of incidence, $\theta=\arctan\left((z-z_{\rm cpa})/r\right)$, versus the distance from the closest point of approach $z-z_{\rm cpa}$ for different source ranges $r$.
The uncertainty in channel locations, the sound speed profile, and the low signal-to-noise ratio make source localization by backpropagation of the amplitude or phase of the recorded signal challenging. Backpropagating the spatial phase gradient $\Delta_{z}\phi$ of the recorded signal allows an accurate estimation of the location along the cable where the source has its closest point of approach. Consider a two-dimensional grid in time $t_{0}\equiv t$ and space $z_{0}\equiv z$ with potential spatiotemporal source locations $(t_{0},z_{0})$, i.e., each point on the grid represents a potential apex of a wavefront as shown in Fig. 12. Assuming plane wave propagation at $\theta=90^{\circ}$, a reasonable approximation for $z-z_{\rm cpa}>10$ km as shown in Fig. 13, summing the data points over all channels $z$ (that is channels 10000–20000) along the wavefront delays $\left|z-z_{0}\right|/c_{a})$ populates the $(z_{0},t_{0})$ grid as,
@@
The two-dimensional backpropagation result normalized to the maximum absolute value $\mathbf{B}/\max(\left|\mathbf{B}\right|)$ is shown in Fig. 14(a). Since the spatial phase gradient changes sign symmetrically around the wavefront’s normal incidence, it will sum up most destructively at the closest point of approach. Figure 14(b) shows the spatial covariance matrix $\mathbf{C}_{b}=\mathbf{B}\mathbf{B}^{T}$ of the two-dimensional backpropagation output $\mathbf{B}$ in Fig. 14(a). The closest point of approach is identified by locating the minimum of the diagonal of the spatial covariance matrix, $z_{\rm cpa}=\arg\min\left(\text{diag}(\mathbf{C}_{b})\right)$; see Fig. 14(c).
Figure 15: (color online) High-pass filtered (cut-off frequency 15 Hz) data featuring a broadband signal attributed to a passing ship. (a) Recordings at successive channels at distance $z$ from the interrogator along the north cable optical fiber. (b) Data representation in the f-k domain.
Figure 16: Estimates of (a) the frequency $f_{a}$ in Eq. (35) and (b) the angle of incidence $\theta$ in Eq. (36) of the acoustic wavefield featuring ship noise sensed by the optical fiber cable.
@@ IV.2 Ship noise
Dataset B comprises acoustic noise data emitted by a cargo ship, collected along the transmit fiber of the north cable on March 11, 2021. The association of the observed signal with ship noise is based on Automatic Identification System (AIS) datawilcock2023distributed indicating that the vessel was located at latitude $45.2600^{\circ}$ N and longitude $124.5477^{\circ}$ W at the time of the dataset B recording, traveling at a speed of 12.6 knots. The duration of the recording is 30 s sampled at 500 Hz. The channel spacing is $L_{h}=2.0419$ m and the gauge length is set to $L_{g}=15L_{h}\approx L_{p}\approx 30$ m.
Similar to Fig. 9, Fig. 15 shows the time-space and f-k representation of the signal recorded over 3000 channels (channels 22500–25500) after de-trending, transforming the measurements to strain units and high-pass filtering at a cut-off frequency of 15 Hz. The recorded signal in dataset B indicates omnidirectional excitation in the frequency range 40–60 Hz, which is attributed to ship noise.wilcock2023distributed The notches along the wavenumber dimension occur at multiples of $k=2\pi/L_{g}=2\pi/30\approx\pm 0.2$ rad/m. Figure 15(a) shows that the amplitude of the ship noise oscillates in time with a period of 12–14 s, which corresponds to the average period of the ocean swell during the experiment. Hence, the observed temporal amplitude modulation can be attributed to changes of the radiated noise due to the ocean swell.arveson2000radiated
Figure 16 shows (a) the acoustic frequency and (b) the incident angle as estimated through the temporal and spatial gradient of the phase of the complex-valued data, respectively. Finally, Fig. 17 shows the result of backpropagating the spatial phase gradient $\Delta_{z}\phi$ of the recorded signal and the estimation of the closest point of approach, similar to Fig. 14.
Figure 18: (color online) T wave signal from the Mw 5.2 Fox Islands earthquake on November 4, 2021 at 08:57:06 UTC. (a) Map with great circle arrival of T wave, (b) bathymetry of the north cable deployment, (c) waveform and spectrogram of the recording from a channel at 40 km from the interrogator, and (d) time-space plot of the recordings along the cable.
@@ IV.3 Earthquake noise
T waves, i.e., acoustic waves coupled into the ocean at a sloping seafloor interface, such as a seamount or the continental margin, due to seismic body waves from nearby earthquakes,okal2008generation can be observed with underwater DAS cables at large epicentral distances.Ugalde2022 T waves from several earthquakes have been observed in the OOI DAS data. Among these, the Mw 5.2 Fox Islands earthquake, which occurred on November 4, 2021, at 08:57:06 UTC, is one of the largest events recorded during this experiment.shen2024ocean
@@
The T wave is observed over a range of around 30 km along the cable, which can be due to a change of the bathymetry reducing the modal cut-off frequency. The reduction of the T wave energy at approximately 60 km along the cable is attributed to the directional sensitivity of DAS, explained in Sec. II.5, since the T wave signal arrives at a 120 heading, resulting in a horizontal incident angle of 60–70 with the cable; see Fig. 18(a).
@@ V.1 Optical multiplexing
Most DAS applications have used dark fibers, i.e., redundant network fibers not used for communications, to avoid interference between overlapping DAS pulses with existing optical traffic. In subsea networks, most fibers are actively used for telecommunications as the cost and complexity of subsea cable deployment limit the number of installed cables. In contrast, in terrestrial networks dark fibers are more readily available. Consequently, dark subsea fibers are rare, and the high demand for intercontinental communication further limits their availability. Optical multiplexing methods based on frequency division can overcome this limitation.
Ordinary infrared optical communication typically occurs in the C-band (“Communication”), wavelength 1530–1565 nm, divided into channels. The information to be transmitted is optically encoded by a transceiver operating at one channel’s wavelength (or equivalently, frequency), and the output of all transceivers is combined into the same fiber—a technique known as wavelength-division multiplexing (WDM)—to permit simultaneous transmission of independent information. Initially, simultaneous operation of DAS and communication traffic in the same fiber was demonstrated with C-band WDM, dedicating three channels for the DAS.huang2019first ; jia2021experimental ; guerrier2022vibration The disadvantages of this approach are that the communication bandwidth is reduced relative to a fiber without DAS and the peak DAS power must be limited to prevent interference effects.jia2021experimental
Another approach utilizes WDM of C-band communications with DAS operating in the L-band (“Long”), wavelength 1565–-1625 nm, which has only slightly higher attenuation than the C-band but is used much less frequently in communications.wienecke2023new L-band multiplexed DAS has been demonstrated both in laboratory,wienecke2023new and in field experiments.brenne2024non ; shi2024multiplexed
@@ V.2 Measurement range
Commercially available DAS systems have a measurement range up to 171 km.waagaard2021realtime Most subsea cables utilize repeaters at regular intervals every 40–100 km to amplify the optical signal and counteract transmission loss. However, common repeaters are effectively one-way gates preventing any backscattered signal from returning back to the interrogator. Consequently, the DAS measurement range is limited to the fiber length that spans the distance between the interrogator and the successive repeater.
@@ V.3 High-frequency signal observation
While most studies have demonstrated the capabilities of DAS in observing hydro-acoustic signals at frequencies below 100 Hz—such as those produced by earthquakes, baleen whales, and ship noise—the maximum frequency for DAS systems is inherently determined by the pulse repetition rate of DAS recordings, which is limited by the two-way travel time of optical pulses along the cable. For instance, a 50 km cable allows a pulse repetition frequency of $f_{r}=c_{n}/(2L_{f})$=2 kHz resulting in a maximum detectable frequency of $f_{\text{max}}=f_{r}/2=1$ kHz. However, several factors can reduce DAS sensitivity to high-frequency signals:
@@
Despite these challenges, several case studies have demonstrated the potential of DAS for observing acoustic signals above 100 Hz. Specifically, ship noise was observed with DAS up to 120 Hz,Thiem2023ShipNC ; cao2023near whereas a sound source emitting frequencies up to 160 Hz was utilized to localize a DAS cable in shallow water.shen2024high Using a pulse repetition frequency of 2 kHz to interrogate shorter fibers (less than 4 km) with smaller gauge lengths, higher frequency signals from impulsive sources have been recorded with DAS up to 700 Hzdouglass2023distributed , 960 Hztaweesintananon2021distributed and 1 kHz.Buisman2022highfreqship Further reducing the interrogation distance, i.e., for fiber lengths up to 1 km, allows for a higher pulse repetition frequency and smaller gauge lengths (4 m) and, consequently, higher observable frequencies up to 2.5 kHz.rychen2023test ; potter2024first ; saw2025distributed DAS has also been shown to record signals at frequencies up to 33 kHz.Buisman2022highfreq
@@ V.4 Unknown geometry
The exact channel locations in DAS are often unknown, unlike hydrophone arrays in ocean acoustics. Particularly, the precise positioning for most telecom cables is either unavailable or proprietary information. Even when the overall cable location is known, estimating the precise channel locations remains challenging. The exact speed of light in optical fiber cannot be measured in situ, leading to uncertainty in converting fast time to distance. Previous studies have demonstrated that a sound source broadcasting near the surface can localize the cable. shen2024high ; Growe2024cableloc By extension, ships with AIS location data can serve as opportunistic sources to help determine the cable’s location.
@@ VI Ocean applications beyond acoustics
Fiber-optic cables deployed on the seafloor are subject to strain not only due to acoustic waves, but also due to mechanical vibrations from earthquakes, slope movement, ocean waves and bottom currents, as well as being exposed to temperature perturbations. This section presents DAS applications in seismology and oceanography to aid in the identification of nonacoustic signals in DAS measurements.
@@ VI.1 Ocean-bottom seismology
Figure 19: (color online) (a) Ambient seismic noise (Scholte waves) from Dataset A along the OOI Regional Cabled Array north cable between channels 9000–14000 at depth $\sim$200 m (see Fig 8), with a second-order Butterworth filter applied between 0.2–1 Hz. (b) Frequency-wavenumber spectrum of the same data segment. Positive wavenumbers indicate propagation towards the shore. (c), (d) Same representations as (a), (b) but for channels 27500–32500 at depth $\sim$500 m.
Seismic monitoring across the oceans is important for early warning of offshore earthquakes and for improving tomographic models of the deep Earth.kohler2020plan Fiber-optic sensing on telecommunication networks deployed across the oceans offers a promising solution to increase the global seismographic coverage. Specifically, DAS has been used for earthquake detection and localization,lior2021detection ; baba2023observation identification of offshore fault zones from scattered waves,lindsey2019illuminating and quantification of linear and nonlinear earthquake ground motion.viens2022nonlinear Recent tests suggest that seismic monitoring with DAS on submarine cables can reduce the early warning time (up to 10 s) for offshore earthquakes.romanowicz2023seafoam ; yin2023real
The horizontal orientation of fiber-optic cables on the seafloor makes DAS particularly sensitive to shear and Scholte (water-sediment interface) waves.martin2021introduction ; spica2020marine Shear-wave resonances observed with DAS can be used to infer the thickness of submarine sediments,trabattoni2024sediment ; taweesintananon2024near while the frequency-dependent dispersion characteristics of surface waves have been used to infer the depth-dependent shear wave velocity profile.spica2020marine ; cheng2021utilizing ; williams2021scholte
Ocean-bottom DAS observations of seismic noise is seen in the 0.2–1 Hz noise radiating from the coast in f-k spectra from the OOI Regional Cabled Array (Fig. 19). Others have used DAS to characterize microseisms.williams2019distributed ; guerin2022quantifying ; xiao2022locating ; fang2022unraveling